Maras analytics firm, Python and Potato, has the following Cobb-Douglas production function: q KL® where a, ß> 0. Mara can
Answer #1

Answer (a) Minimize : Total lest (Tc) = ukrwl subject to : K P = q (1) Legrange is given by 2= uk + WL + aq -K ) First Order+ = a 9 awa+A -BV - optimal value Sk - = 8vk - +B sy aw -- Chintal Iaw value of L ratal cak (TE); wh+VK - TLE G - లో ఈ - 742* ze =(w3 06 - Total Cost function (C)___ Marginal cost (MC) = d TC ate 1-2-B X q atß 1-2-8 => MCB ( q 276 2009 X+B Marginal